# Hyperfocal subalgebras of hyperfocal abelian Frobenius blocks

### v1

@@
Xueqin Hu1, Kun Zhang2 and Yuanyang Zhou1
@@ 1 Introduction
In the light of Rouquier’s conjecture, some recent work focuses on the investigation of the structure of the block algebra $\mbox{$\mathcal{O}$}Gb$ with some special abelian hyperfocal subgroup. In [33] and [7], the numbers of ordinary irreducible characters and irreducible Brauer characters were obtained for the blocks with cyclic and Klein four hyperfocal subgroup, respectively. In [8], both authors constructed an isotypy, which is a character-theoretic shadow of Rouquier’s conjecture, between the blocks $b$ and $c$ when the hyperfocal subgroup $D$ is cyclic and the quotient group $P/D$ is abelian. These results only reflect the influences of the structure of hyperfocal subgroups on the structure of characters of block algebras. It may be more interesting and meaningful if the algebraic structure of the block algebras with some special abelian hyperfocal subgroups can be determined.
@@
Motivated by this, in this paper, we will focus on the structure of the hyperfocal subalgebra of a block with a special class of hyperfocal subgroups, which includes the two classical cases where the hyperfocal subgroup is a Klein four group and cyclic, respectively. To make this explicit, we need some notation. Let $\mathbb{A}$ be a hyperfocal subalgebra of the block $b$. Let $D_{\delta}$ be a local pointed group of $D$ on the block algebra $\mbox{$\mathcal{O}$}Gb$ and denote by $N_{G}(D_{\gamma})$ its stabilizer under the conjugation action of $G$ (see the paragraph 2.9 below). As it is shown in Proposition 3.7 below, the quotient group $N_{G}(D_{\gamma})/DC_{G}(D)$ is $p$-nilpotent. Its normal $p$-complement is denoted by $E_{\mathfrak{hyp}}$. Since the order of $E_{\mathfrak{hyp}}$ is coprime to $p$, we can view it as a subgroup of the automorphism group $\mathrm{Aut}(D)$ of $D$. Then the block $b$ is called with Frobenius hyperfocal inertial quotient if $E_{\mathfrak{hyp}}$ acts freely on $D-\{1\}$, namely the set of all nontrivial elements of $D$ (see the paragraph 3.8 below). This is equivalent to requiring that the semidirect product $D\rtimes E_{\mathfrak{hyp}}$ is a Frobenius group. Now we can state the main result of this paper as follows.
Theorem 1.1. (see also Theorem 5.14) Keep the notation as above. Suppose that the hyperfocal subgroup $D$ is abelian and the block $b$ is with Frobenius hyperfocal inertial quotient. Then there is a stable equivalence of Morita type between the hyperfocal subalgebra $\mathbb{A}$ and the group algebra $\mbox{$\mathcal{O}$}(D\rtimes E_{\mathfrak{hyp}})$.
Throughout this paper, a block with abelian hyperfocal subgroup and Frobenius hyperfocal inertial quotient is called a hyperfocal abelian Frobenius block for simplicity. By definition, the hyperfocal abelian Frobenius block is an analogous version of the block with abelian defect group and Frobenius inertial quotient, which is called an abelian Frobenius block here for simplicity, at the hyperfocal level. So the main theorem above can be viewed as a hyperfocal version of [17, Theorem 10.5.1], which is due to Puig (see [23, 6.8]). Just like abelian Frobenius blocks including the blocks with Klein four defect groups and cyclic defect groups, the hyperfocal abelian Frobenius blocks include the blocks with Klein four hyperfocal subgroups and cyclic hyperfocal subgroups (see Remark 3.9 below). Thus we can apply this main theorem above to the blocks with Klein four hyperfocal subgroups and cyclic hyperfocal subgroups and obtain the structures of the corresponding hyperfocal subalgebras as follows, which are similar to the structures of the blocks with Klein four defect groups and cyclic defect groups, respectively.
@@
Theorem 1.3. (see also Theorem 6.20) Assume that the hyperfocal subgroup $D$ is a cyclic group. Then the hyperfocal subalgebra $\mathbb{A}$ is Rickard equivalent to the group algebra $\mbox{$\mathcal{O}$}(D\rtimes E_{\mathfrak{hyp}})$.
Theorem 1.3 is a generalization of [9, Theorem 1.2] since we don’t need the defect group being abelian and then gives a complete description of the structure of the hyperfocal subalgebra with cyclic hyperfocal subgroup. This result also gives a positive answer to a question whether the hyperfocal subalgebra with nontrivial cyclic hyperfocal subgroup is a Brauer tree algebra proposed by Professor Linckelmann some years ago.
Furthermore, by using Clifford theoretic relationship between the representation theory of the source algebra and its hyperfocal subalgebra, we can also obtain some information about the characters of the blocks with Klein four hyperfocal subgroups and cyclic hyperfocal subgroups in terms of source algebras. We refer to Propositions 6.15 and 6.34 for detailed descriptions of these structures. A key step towards these descriptions is to determine the relationship between the dimenisons of simple $\mathcal{K}\mathop{\otimes}\limits_{\mathcal{O}}\mathbb{A}$-modules and the prime number $p$. At the same time, this relates a conjecture, denoted by KLN cojecture for simplicity, proposed by Kessar and Linckelmann and Navarro in [13]. We recall this conjecture in the following.
@@
KLN Conjecture. ([13, Conjecture 1.2]) With the notation above, assume that $\mathcal{K}$ is a splitting field of $\mathbb{A}$. Then $\mathrm{Irr}_{\mathcal{K}}(\mathbb{A})=\mathrm{Irr}_{\mathcal{K},0}(\mathbb{A})$ if and only if the hyperfocal subgroup $D$ is abelian.
In [13], the KLN conjecture is viewed as a ‘hyperfocal height zero’ version of Brauer’s height zero conjecture. In this paper, we can verify the forward direction of the KLN conjecture for the cases where the hyperfocal subgroup $D$ is a Klein four group and cyclic.
@@
Proposition 1.5. (see also Proposition 6.28) When the hyperfocal subgroup $D$ is cyclic, then $\mathrm{Irr}_{\mathcal{K}}(\mathbb{A})=\mathrm{Irr}_{\mathcal{K},0}(\mathbb{A})$.
@@
Now let us illustrate the construction of the stable equivalence of Morita type in Theorem 1.1 briefly. This construction is similar to the one in [17, Theorem 10.5.1] for abelian Frobenius blocks. The main ingredients used for the construction in [17, Theorem 10.5.1] are the descriptions of local structures of an abelian Frobenius block and Puig’s Gluing Theorem for indecomposable endopermutation modules of abelian $p$-groups. Let us make it more explicit. For an abelian Frobenius block with defect group $P$, its Brauer correspondents in the centralizers of nontrivial subgroups of $P$ are nilpotent in the sense of [2]. By the structure theorem of the source algebras of nilpotent blocks, due to Puig [22], these local structures can yield a family of indecomposable endopermutation modules (see the paragraph 2.14 below) of quotient groups of the abelian $p$-group $P$ satisfying some compatibility assumptions. Then by Puig’s Gluing Theorem, these endopermutation modules can be glued together to an indecomposable endopermutation module of $P$, which plays an essential role in the construction of the stable equivalence of Morita type for abelian Frobenius blocks.
@@
Analogously, this method can be carried over to the hyperfocal algebra of the hyperfocal abelian Frobenius block. Instead of describing the structures of the Brauer correspondents in the centralizers of nontrivial subgroups of the defect group $P$, we need to determine the structures of the Brauer correspondents in the centralizers of nontrivial subgroups of the hyperfocal subgroup $D$, which are shown to be also nilpotent by Proposition 3.12 below. Naturally, the next step should be applying the structure theorem of the source algebras of nilpotent blocks to these blocks to yield a family of indecomposable endopermutation modules of quotient groups of $D$. However, there is an obstruction to this application because the Brauer quotients of the source algebra at nontrivial subgroups of $D$ may not be a source algebra of the block of the corresponding centralizer when the defect group $P$ is nonabelian. Due to this obstruction, instead of considering the source algebra and the hyperfocal subalgebra, we will focus on a so-called embedded algebra $\mathbb{A}_{\delta}$ of the pointed group $D_{\delta}$ of the hyperfocal subalgebra $\mathbb{A}$ (see the paragraph 2.10 below), which can be viewed as a subalgebra (not unitary in general) of the hyperfocal subalgebra $\mathbb{A}$. The embedded algebra $\mathbb{A}_{\delta}$ has some properties similar to the source algebra, such as the existence of a Morita equivalence between $\mathbb{A}_{\delta}$ and the hyperfocal subalgebra $\mathbb{A}$ (see Corollary 3.4 below) and an $\mbox{$\mathcal{O}$}(D\times D)$-module structure similar to the $\mbox{$\mathcal{O}$}(P\times P)$-module structure of the source algebra (see Proposition 4.6 below). These properties will play essential roles in the descriptions of the structures of the Brauer quotients of $\mathbb{A}_{\delta}$ at nontrivial subgroups of $D$ and eventually the construction of the stable equivalence of Morita type we need.
@@
The paper is organized as follows. In Section $2$, some necessary notation and definitions are collected. In Section $3$, the definition of the hyperfocal abelian Frobenius block is introduced after we show that the quotient group $N_{G}(D_{\delta})/DC_{G}(D)$ is $p$-nilpotent. Then the local structures of this block associated with nontrivial subgroups of the hyperfocal subgroup are determined. Also, some basic properties of the embedded algebra $\mathbb{A}_{\delta}$ are proved in this section. Section $4$ is devoted to determining the structures of the embedded algebra $\mathbb{A}_{\delta}$, including its structure as an $\mbox{$\mathcal{O}$}(D\times D)$-module and the algebraic structure of its Brauer quotient at every nontrivial subgroup of the hyperfocal subgroup, which are essential to the construction of a stable equivalence of Morita type done in Section $5$. Besides this construction, in Section $5$, we also investigate an equivariant property of the bimodule inducing this stable equivalence, which is necessary when we want to obtain the information about the source algebra. In Section $6$, the structures of the hyperfocal algebras of blocks with Klein four hyperfocal subgroups and cyclic hyperfocal subgroups are determined as applications of the main result obtained in Section $5$. Some information about characters of these two blocks is also obtained in Section $6$.
@@ 2 Preliminaries
Let $S$ be a finite set. We denote by $|S|$ the cardinality of $S$. Let $H$ be a finite group and $\mathfrak{b}$ a block of $H$ over $\mathcal{O}$. Denote by $\mathrm{Irr}(H)$ and $\mathrm{Irr}(\mathfrak{b})$ the set of ordinary irreducible characters of $H$ and the set of ordinary irreducible characters of $\mathfrak{b}$, respectively. Similary, we denote by $\mathrm{IBr}(H)$ and $\mathrm{IBr}(\mathfrak{b})$ the set of modular irreducible characters of $H$ and the set of modular irreducible characters of $\mathfrak{b}$, respectively. If no confusion arises, we denote by $1_{H}$ the trivial character of $H$. Let $H_{1}$ and $H_{2}$ be two subgroups of $H$. For any $\xi\in\mathrm{Irr}(H_{1})$, we denote by $\mathrm{Ind}_{H_{1}}^{H}(\xi)$ the induced character of $\xi$. We use the similar notation for the induced module. Analogously, we denote the restriction by $\mathrm{Res}_{H_{1}}^{H}$. For any $\mathcal{O}H_{1}$-module $M$ and any $h\in H$, we define the twisted $\mathcal{O}(hH_{1}h^{-1})$-module $_{h}M$ by setting $_{h}M=M$ as an $\mathcal{O}$-module, with $hh_{1}h^{-1}\in hH_{1}h^{-1}$ acting on $m\in{{}_{h}M}$ as $h_{1}$ on $m\in M$. Furthermore, if $M$ is an $\mathcal{O}H_{1}$-$\mathcal{O}H_{2}$-bimodule, then by preserving the right $\mathcal{O}H_{2}$-module structure, $_{h}M$ is an $\mathcal{O}(hH_{1}h^{-1})$-$\mathcal{O}H_{2}$-bimodule.
@@
Throughout this paper, an $\mathcal{O}$-algebra always satisfies the property that it is free and has finite rank as an $\mathcal{O}$-module. An $\mathcal{O}$-algebra $C$ is called an indecomposable algebra if it is indecomposable as a $C$-$C$-bimodule, which is equivalent to saying that its center $Z(C)$ is local. Let $\tilde{C}$ be a subalgebra of $C$. We denote by $\mathrm{Res}_{\tilde{C}}(M)$ the restriction of $M$ to $\tilde{C}$ for any left $C$-module $M$. Denote by $1_{C}$ (or 1 if no confusion arises), $J(C)$ and $C^{*}$ the unit element of $C$, the Jacobson radical of $C$, the set of invertible elements of $C$, respectively. We denote by $C^{\circ}$ the opposite algebra of $C$. Let $C^{\prime}$ be another $\mathcal{O}$-algebra. Let $\varphi$ be an isomorphism from $C$ to $C^{\prime}$ and $M$ a left $C$-module. We use ${{}_{\varphi}}M$ to denote the left $C^{\prime}$-module with being equal to $M$ as $\mathcal{O}$-modules and the left action of $c^{\prime}\in C^{\prime}$ on $_{\varphi}C$ being equal to the left action of $\varphi^{-1}(c^{\prime})$ on $M$. For any $CC^{\prime}-$bimodule $M$, we can define $(x\mathop{\otimes}\limits_{\mathcal{O}}x^{\prime})m=xmx^{\prime}$ for any $x\in C$, $x^{\prime}\in(C^{\prime})^{\circ}$ and $m\in M$. In this way, it can be viewed as a $C\mathop{\otimes}\limits_{\mathcal{O}}(C^{\prime})^{\circ}$-module and vice versa. Throughout this paper, we identify $CC^{\prime}-$bimodules with $C\mathop{\otimes}\limits_{\mathcal{O}}(C^{\prime})^{\circ}$-modules through this way and vice versa.
@@
We denote the Heller operator of $C$ by $\Omega_{C}$ for an $\mathcal{O}$-algebra $C$. This operator can induce an equivalence on the $\mathcal{O}$-stable category of $\mathrm{mod}(C)$, of which the definition we recall as follows. Recall that a $C$-module $U$ is called relative $\mathcal{O}$-projective, if it is isomorphic to a direct summand of $C\mathop{\otimes}\limits_{\mathcal{O}}V$ for some $\mathcal{O}$-module $V$ (see [16, Definition 2.6.11]). Note that if the $C$-module $U$ is relative $\mathcal{O}$-projective and free of finite rank, then $U$ is a projective $C$-module. Denote by $\mathrm{mod}(C)$ the category of finitely generated left $C$-modules. The quotient category of $\mathrm{mod}(C)$ by the subcategory of relatively $\mathcal{O}$-projective modules is called $\mathcal{O}$-stable category of $\mathrm{mod}(C)$ (see [14, 11.1]). We denote it by $\overline{\mathrm{mod}}(C)$.
@@
Keep the notation above. Denote $\mathrm{Hom}_{\mathcal{O}}(C,\mathcal{O})$ by $C^{*}$. It is the dual module of the $C$-$C$-bimodule $C$. The $\mathcal{O}$-algebra $C$ is called symmetric over $\mathcal{O}$ if $C$ is finitely generated projective as an $\mathcal{O}$-module and $C\cong C^{*}$ as $C$-$C$-bimodules (see [16, Definition 2.11.1]). Take two symmetric $\mathcal{O}$-algebras $C$ and $C^{\prime}$. Let $M_{\cdot}$ be a bounded complex of finitely generated $CC^{\prime}-$bimodules which are projective as left $C$-modules and right $C^{\prime}$-modules. Set $M_{\cdot}^{*}$ to be the dual $\mathrm{Hom}_{\mathcal{O}}(M_{\cdot},\mathcal{O})$ of the complex $M_{\cdot}$. We say that the complex $M_{\cdot}$ induces a Rickard equivalence between $C^{\prime}$ and $C$, if there is a $CC^{\prime}-$bimodule $F$ and a $^{\prime}-CC$-bimodule $F^{\prime}$ such that $F$ and $F^{\prime}$ are homotopy equivalent to 0 as complexes of bimodules, and if $M_{\cdot}\mathop{\otimes}\limits_{C^{\prime}}M_{\cdot}^{*}\cong C\oplus F$ as complexes of $C$-$C$-bimodules and $M_{\cdot}^{*}\mathop{\otimes}\limits_{C}M_{\cdot}\cong C^{\prime}\oplus F^{\prime}$ as complexes of $^{\prime}-CC^{\prime}$-bimodules (see [27, Definitions 2.2.1]). The complex $M_{\cdot}$ is called a Rickard complex. It is clear that if $C$ and $C^{\prime}$ are Rickard equivalent, then they are derived equivalent, namely their bounded derived categories $\mathcal{D}^{b}(C)$ and $\mathcal{D}^{b}(C^{\prime})$ are equivalent to each other as triangulated categories.
In the following two paragraphs, we will collect some notation from [12, §2-§3] which will be used later in the descriptions of characters of the source algebras and the hyperfocal subalgebras in Section $6$. Throughout the following $2$ paragraphs, we fix a symmetric $\mathcal{O}$-algebra $C$ such that $\mathcal{K}\mathop{\otimes}_{\mathcal{O}}C$ is a split semi-simple $\mathcal{K}$-algebra, which means that $\mathcal{K}\mathop{\otimes}_{\mathcal{O}}C$ is isomorphic to a direct product of matrix algebras over $\mathcal{K}$.

### v2

@@
Xueqin Hu1, [^1] E-mail address: hxq@mail.ccnu.edu.cn ,   Kun Zhang2 and Yuanyang Zhou1
@@ 1 Introduction
In the light of Rouquier’s conjecture, some recent work focuses on understanding the structure of the block algebra $\mbox{$\mathcal{O}$}Gb$ with some special abelian hyperfocal subgroup. In [33] and [7], the numbers of ordinary irreducible characters and irreducible Brauer characters were determined for blocks with a cyclic or a Klein four hyperfocal subgroup, respectively. In [8], the first and third authors constructed an isotypy, which is a character-theoretic shadow of Rouquier’s conjecture, between the blocks $b$ and $c$ when the hyperfocal subgroup $D$ is cyclic and the quotient group $P/D$ is abelian. These results reveal the influences of the hyperfocal subgroup on the character structure of block algebras. It may be more interesting and meaningful if the algebra structure of the block algebras with some special abelian hyperfocal subgroups can be determined.
@@
Motivated by this, in this paper, we will focus on the structure of the hyperfocal subalgebra of a block with a special class of hyperfocal subgroups, which includes the two classical cases where the hyperfocal subgroup is a Klein four group or cyclic, respectively. To make this explicit, we introduce some notation. Let $\mathbb{A}$ be a hyperfocal subalgebra of the block $b$. Let $D_{\delta}$ be a local pointed group of $D$ on the block algebra $\mbox{$\mathcal{O}$}Gb$ and denote by $N_{G}(D_{\delta})$ its stabilizer under the conjugation action of $G$ (see the paragraphs 2.9-2.11 below). As shown in Proposition 3.7 below, the quotient group $N_{G}(D_{\delta})/DC_{G}(D)$ is $p$-nilpotent. Its normal $p$-complement is denoted by $E_{\mathfrak{hyp}}$. Since the order of $E_{\mathfrak{hyp}}$ is coprime to $p$, we can view it as a subgroup of the automorphism group $\mathrm{Aut}(D)$ of $D$. Then the block $b$ is called with a Frobenius hyperfocal inertial quotient if $E_{\mathfrak{hyp}}$ acts freely on $D-\{1\}$, namely the set of all nontrivial elements of $D$ (see the paragraph 3.8 below). This is equivalent to requiring that the semidirect product $D\rtimes E_{\mathfrak{hyp}}$ is a Frobenius group. Now we can state the main result of this paper as follows.
Theorem 1.1. (see also Theorem 5.14) Keep the notation as above. Suppose that the hyperfocal subgroup $D$ is abelian and the block $b$ is with a Frobenius hyperfocal inertial quotient. Then there is a stable equivalence of Morita type between the hyperfocal subalgebra $\mathbb{A}$ and the group algebra $\mbox{$\mathcal{O}$}(D\rtimes E_{\mathfrak{hyp}})$.
Throughout this paper, a block with an abelian hyperfocal subgroup and a Frobenius hyperfocal inertial quotient is called a hyperfocal abelian Frobenius block for short. By definition, such block is the hyperfocal analogue of the block with an abelian defect group and a Frobenius inertial quotient, which is called an abelian Frobenius block here for simplicity. Thus, the main theorem above can be viewed as a hyperfocal version of [16, Theorem 10.5.1], which is due to Puig (see [23, 6.8]). Just as abelian Frobenius blocks include the blocks with Klein four defect groups or cyclic defect groups, the hyperfocal abelian Frobenius blocks include those with Klein four hyperfocal subgroups or cyclic hyperfocal subgroups (see Remark 3.9 below). Thus we can apply this main theorem to the blocks with Klein four hyperfocal subgroups or cyclic hyperfocal subgroups and obtain the structure of the corresponding hyperfocal subalgebras as follows, which are similar to the structure of the blocks with Klein four defect groups or cyclic defect groups, respectively.
@@
Theorem 1.3. (see also Theorem 6.22) Assume that the hyperfocal subgroup $D$ is a nontrivial cyclic group. Then the hyperfocal subalgebra $\mathbb{A}$ is Rickard equivalent to the group algebra $\mbox{$\mathcal{O}$}(D\rtimes E_{\mathfrak{hyp}})$.
Theorem 1.3 is a generalization of [9, Theorem 1.2], as we no longer require the defect group to be abelian, and thus provides a complete description of the hyperfocal subalgebra with a nontrivial cyclic hyperfocal subgroup. This result also answers a question proposed by Professor Linckelmann some years ago, asking whether such a hyperfocal subalgebra is a Brauer tree algebra.
Furthermore, by using the Clifford theoretic relationship between the representation theory of the source algebra and its hyperfocal subalgebra, we can also obtain some information about the characters of the source algebra of a block with a Klein four hyperfocal subgroup or a cyclic hyperfocal subgroup. We refer to Propositions 6.17 and 6.37 for detailed descriptions of these structures. A key step towards these descriptions is to determine the relationship between the dimenisons of simple $\mathcal{K}\mathop{\otimes}\limits_{\mathcal{O}}\mathbb{A}$-modules and the prime $p$. This is closely related to a conjecture proposed by Kessar and Linckelmann and Navarro in [12], which we will refer to as the KLN conjecture. We recall this conjecture in the following.
@@
KLN Conjecture. ([12, Conjecture 1.2]) With the notation above, assume that $\mathcal{K}$ is a splitting field of $\mathbb{A}$. Then $\mathrm{Irr}_{\mathcal{K}}(\mathbb{A})=\mathrm{Irr}_{\mathcal{K},0}(\mathbb{A})$ if and only if the hyperfocal subgroup $D$ is abelian.
In [12], the KLN conjecture is viewed as a ‘hyperfocal height zero’ version of Brauer’s height zero conjecture. In this paper, we can verify the forward direction of the KLN conjecture in the cases where the hyperfocal subgroup $D$ is a Klein four group or a nontrivial cyclic group.
@@
Proposition 1.5. (see also Proposition 6.30) When the hyperfocal subgroup $D$ is nontrivial cyclic, then $\mathrm{Irr}_{\mathcal{K}}(\mathbb{A})=\mathrm{Irr}_{\mathcal{K},0}(\mathbb{A})$.
@@
Now let us now briefly outline the construction of the stable equivalence of Morita type stated in Theorem 1.1. This construction is analogous to that in [16, Theorem 10.5.1] for abelian Frobenius blocks. The main ingredients there are the descriptions of local structures of an abelian Frobenius block and Puig’s Gluing Theorem for indecomposable endopermutation modules of abelian $p$-groups. Let us make it more explicit. For an abelian Frobenius block with defect group $P$, its Brauer correspondents in the centralizers of nontrivial subgroups of $P$ are nilpotent in the sense of [2]. By the structure theorem of the source algebras of nilpotent blocks, due to Puig [22], these local structures can yield a family of indecomposable endopermutation modules (see the paragraph 2.14 below) of quotient groups of the abelian $p$-group $P$ satisfying certain compatibility conditions. Then by Puig’s Gluing Theorem, these endopermutation modules can be glued together to an indecomposable endopermutation module of $P$, which plays a crucial role in the construction of the stable equivalence of Morita type for abelian Frobenius blocks.
@@
We adapt this approach to the hyperfocal subalgebra of a hyperfocal abelian Frobenius block. Instead of analyzing the Brauer correspondents in the centralizers of nontrivial subgroups of the defect group $P$, we need to understand the Brauer correspondents in the centralizers of nontrivial subgroups of the hyperfocal subgroup $D$, which are shown to be also nilpotent by Proposition 3.12 below. Naturally, the next step should be applying the structure theorem of the source algebras of nilpotent blocks to these blocks to yield a family of indecomposable endopermutation modules of quotient groups of $D$. However, an obstruction arises when the defect group $P$ is nonabelian because the Brauer quotients of the source algebra at nontrivial subgroups of $D$ may not be the source algebras of the blocks of the corresponding centralizers. To circumvent this, instead of considering the source algebra and the hyperfocal subalgebra, we will focus on a so-called embedded algebra $\mathbb{A}_{\delta}$ of the pointed group $D_{\delta}$ of the hyperfocal subalgebra $\mathbb{A}$ (see the paragraph 2.9 below), which can be viewed as a subalgebra (not unitary in general) of the hyperfocal subalgebra $\mathbb{A}$. The embedded algebra $\mathbb{A}_{\delta}$ shares several properties with the source algebra, such as the existence of a Morita equivalence between $\mathbb{A}_{\delta}$ and the hyperfocal subalgebra $\mathbb{A}$ (see Corollary 3.4 below) and an $\mbox{$\mathcal{O}$}(D\times D)$-module structure similar to the $\mbox{$\mathcal{O}$}(P\times P)$-module structure of the source algebra (see Proposition 4.6 below). These properties are essential for describing the Brauer quotients of $\mathbb{A}_{\delta}$ at nontrivial subgroups of $D$ and eventually for constructing the desired stable equivalence of Morita type.
@@
The paper is organized as follows. In Section $2$, some necessary notation and definitions are collected. In Section $3$, the definition of the hyperfocal abelian Frobenius block is introduced after we show that the quotient group $N_{G}(D_{\delta})/DC_{G}(D)$ is $p$-nilpotent. Then the local structures of such block associated with nontrivial subgroups of the hyperfocal subgroup are determined. Also, some basic properties of the embedded algebra $\mathbb{A}_{\delta}$ are established in this section. Section $4$ is devoted to detailed study of the structures of the embedded algebra $\mathbb{A}_{\delta}$, including its $\mbox{$\mathcal{O}$}(D\times D)$-module structure and the algebra structure of its Brauer quotient at every nontrivial subgroup of the hyperfocal subgroup $D$. These results are crucial for the construction of a stable equivalence of Morita type in Section $5$. In addition to this construction, we also investigate in Section $5$ an equivariant property of the bimodule inducing this stable equivalence, which is necessary to describe the structure of the source algebra. Finally, in Section $6$, we apply the main result of Section 5 to blocks with a Klein four hyperfocal subgroup or a cyclic hyperfocal subgroup, determining the structures of their hyperfocal subalgebras and obtaining some character-theoretic information of their source algebras.
@@ 2 Preliminaries
Let $S$ be a finite set. We denote by $|S|$ the cardinality of $S$. Let $H$ be a finite group and $\mathfrak{b}$ a block of $H$ over $\mathcal{O}$. Denote by $\mathrm{Irr}(H)$ and $\mathrm{Irr}(\mathfrak{b})$ the set of ordinary irreducible characters of $H$ and the set of ordinary irreducible characters of $\mathfrak{b}$, respectively. Similary, we denote by $\mathrm{IBr}(H)$ and $\mathrm{IBr}(\mathfrak{b})$ the set of modular irreducible characters of $H$ and the set of modular irreducible characters of $\mathfrak{b}$, respectively. Let $H_{1}$ and $H_{2}$ be two subgroups of $H$. For any $\xi\in\mathrm{Irr}(H_{1})$, we denote by $\mathrm{Ind}_{H_{1}}^{H}(\xi)$ the induced character of $\xi$. We use the similar notation for the induced module. Analogously, we denote the restriction by $\mathrm{Res}_{H_{1}}^{H}$. For any $\mathcal{O}H_{1}$-module $M$ and any $h\in H$, we define the twisted $\mathcal{O}(hH_{1}h^{-1})$-module $_{h}M$ by setting $_{h}M=M$ as an $\mathcal{O}$-module, with $hh_{1}h^{-1}\in hH_{1}h^{-1}$ acting on $m\in{{}_{h}M}$ as $h_{1}$ on $m\in M$. Furthermore, if $M$ is an $\mathcal{O}H_{1}$-$\mathcal{O}H_{2}$-bimodule, then by preserving the right $\mathcal{O}H_{2}$-module structure, $_{h}M$ is an $\mathcal{O}(hH_{1}h^{-1})$-$\mathcal{O}H_{2}$-bimodule. At the same time, $M$ can be viewed as an $\mbox{$\mathcal{O}$}(H_{1}\times H_{2})$-module with the action of $(h_{1},h_{2})\in H_{1}\times H_{2}$ on $m\in M$ defined as $h_{1}mh_{2}^{-1}$ and vice versa.
@@
Throughout this paper, an $\mathcal{O}$-algebra always satisfies the properties that it is free and has finite rank as an $\mathcal{O}$-module. An $\mathcal{O}$-algebra $C$ is called an indecomposable algebra if it is indecomposable as a $C$-$C$-bimodule, which is equivalent to saying that its center $Z(C)$ is local. Let $\tilde{C}$ be a subalgebra of $C$. We denote by $\mathrm{Res}_{\tilde{C}}(M)$ the restriction of $M$ to $\tilde{C}$ for any left $C$-module $M$. Denote by $1_{C}$ (or 1 if no confusion arises), $J(C)$ and $C^{\times}$ the unit element of $C$, the Jacobson radical of $C$, the set of invertible elements of $C$, respectively. We denote by $C^{\circ}$ the opposite algebra of $C$. Let $C^{\prime}$ be another $\mathcal{O}$-algebra. Let $\varphi$ be a homomorphism from $C$ to $C^{\prime}$ and $M^{\prime}$ a left $C^{\prime}$-module. We use ${{}_{\varphi}}M^{\prime}$ to denote the left $C$-module with being equal to $M^{\prime}$ as $\mathcal{O}$-modules and the left action of $c\in C$ on $_{\varphi}M^{\prime}$ being equal to the left action of $\varphi(c)$ on $M^{\prime}$. For any $CC^{\prime}-$bimodule $\mathbb{M}$, we can define $(x\mathop{\otimes}\limits_{\mathcal{O}}x^{\prime}){\rm m}=x{\rm m}x^{\prime}$ for any $x\in C$, any $x^{\prime}\in(C^{\prime})^{\circ}$ and any ${\rm m}\in\mathbb{M}$. In this way, it can be viewed as a $C\mathop{\otimes}\limits_{\mathcal{O}}(C^{\prime})^{\circ}$-module and vice versa. Throughout this paper, we identify $CC^{\prime}-$bimodules with $C\mathop{\otimes}\limits_{\mathcal{O}}(C^{\prime})^{\circ}$-modules through this way and vice versa.
@@
We denote the Heller operator of $C$ by $\Omega_{C}$ for an $\mathcal{O}$-algebra $C$. This operator can induce an equivalence on the $\mathcal{O}$-stable category of $\mathrm{mod}(C)$, of which the definition we recall as follows. Recall that a $C$-module $U$ is called relative $\mathcal{O}$-projective, if it is isomorphic to a direct summand of $C\mathop{\otimes}\limits_{\mathcal{O}}V$ for some $\mathcal{O}$-module $V$ (see [15, Definition 2.6.11]). Note that if the $C$-module $U$ is relative $\mathcal{O}$-projective and free of finite rank, then $U$ is a projective $C$-module. Denote by $\mathrm{mod}(C)$ the category of finitely generated left $C$-modules. The quotient category of $\mathrm{mod}(C)$ by the subcategory of relatively $\mathcal{O}$-projective modules is called the $\mathcal{O}$-stable category of $\mathrm{mod}(C)$ (see [13, 11.1]). We denote it by $\overline{\mathrm{mod}}(C)$.
@@
Keep the notation above. For any $C$-module $M$, we denote $\mathrm{Hom}_{\mathcal{O}}(M,\mathcal{O})$ by $M^{*}$. It is the dual module of the $C$-module $M$. The $\mathcal{O}$-algebra $C$ is called symmetric over $\mathcal{O}$ if $C$ is a finitely generated projective $\mathcal{O}$-module and $C\cong C^{*}$ as $C$-$C$-bimodules (see [15, Definition 2.11.1]). Take two symmetric $\mathcal{O}$-algebras $C$ and $C^{\prime}$. Let $M_{\cdot}$ be a bounded complex of finitely generated $CC^{\prime}-$bimodules which are projective as left $C$-modules and right $C^{\prime}$-modules. Set $M_{\cdot}^{*}$ to be the dual $\mathrm{Hom}_{\mathcal{O}}(M_{\cdot},\mathcal{O})$ of the complex $M_{\cdot}$. We say that the complex $M_{\cdot}$ induces a Rickard equivalence between $C^{\prime}$ and $C$, if there is a $CC^{\prime}-$bimodule $F$ and a $^{\prime}-CC$-bimodule $F^{\prime}$ such that $F$ and $F^{\prime}$ are homotopy equivalent to 0 as complexes of bimodules, and if $M_{\cdot}\mathop{\otimes}\limits_{C^{\prime}}M_{\cdot}^{*}\cong C\oplus F$ as complexes of $C$-$C$-bimodules and $M_{\cdot}^{*}\mathop{\otimes}\limits_{C}M_{\cdot}\cong C^{\prime}\oplus F^{\prime}$ as complexes of $^{\prime}-CC^{\prime}$-bimodules (see [27, Definitions 2.2.1]). The complex $M_{\cdot}$ is called a Rickard complex. It is clear that if $C$ and $C^{\prime}$ are Rickard equivalent, then they are derived equivalent, namely their bounded derived categories $\mathcal{D}^{b}(C)$ and $\mathcal{D}^{b}(C^{\prime})$ are equivalent to each other as triangulated categories.
In the following two paragraphs, we will collect some notation from [11, §2-§3] which will be used later in the descriptions of characters of the source algebras and the hyperfocal subalgebras in Section $6$. Throughout the following $2$ paragraphs, we fix a symmetric $\mathcal{O}$-algebra $C$ such that $\mathcal{K}\mathop{\otimes}_{\mathcal{O}}C$ is a split semi-simple $\mathcal{K}$-algebra, which means that $\mathcal{K}\mathop{\otimes}_{\mathcal{O}}C$ is isomorphic to a direct product of matrix algebras over $\mathcal{K}$.